OpenStudy (anonymous):

Find the exact value of cos(arcsin(one fourth)). For full credit, explain your reasoning.

OpenStudy (dumbcow):

\[\theta = \sin^{-1} (1/4)\] which means \[\sin \theta = \frac{1}{4}\] you have to find cos theta pythagorean identity is: \[\sin^{2} \theta + \cos^{2} \theta = 1\] \[\frac{1}{16} + \cos^{2} \theta = 1\] \[\cos \theta = \pm \sqrt{1-\frac{1}{16}}\]

OpenStudy (anonymous):

@dumbcow doesn't the answer need to be something like pi/4 or 2pi/3 ?

OpenStudy (dumbcow):

hmm no because they are not asking for the angle but the cos of the angle "arcsin 1/4"

OpenStudy (anonymous):

@dumbcow well all of my other questions have had an answer like that.so i just thought this one needed an answer like that also.

OpenStudy (dumbcow):

sin(x) gives a ratio between -1 and 1 arcsin(x) gives an angle (usually in terms of pi) so the type of answer depends on the question and what you are asked to find

OpenStudy (anonymous):

the question is "Find the exact value of cos(arcsin(one fourth)). For full credit, explain your reasoning." but the multiple choice questions before all had answers like pi/4 or pi/3 and they said the same thing

OpenStudy (dumbcow):

are you sure it was exactly the same thing here is solution from wolfram http://www.wolframalpha.com/input/?i=cos%28arcsin%281%2F4%29%29

OpenStudy (anonymous):



OpenStudy (anonymous):

there is an angle with \(\sin(\theta)=\frac{1}{4}\) adjacent side is , via pythagoras, \(\sqrt{4^2-1^2}=\sqrt{15}\)

OpenStudy (anonymous):

@dumbcow yes i am sure, do you think you can help me out with the answer that relates to that?

OpenStudy (anonymous):

this tells you \[\cos(\theta)=\frac{\sqrt{15}}{4}\]

OpenStudy (dumbcow):

triangle approach works great too thanks @satellite73

OpenStudy (anonymous):

either way, i think the picture is easiest, but it all comes down to pythagoras in any case amazing how much mileage comes out of that little theorem

OpenStudy (anonymous):

alright, thank. @satellite73 @dumbcow

OpenStudy (dumbcow):

yw, if your multiple choice answers all are angles then your teacher had a typo :{